OpenStudy (anonymous):

x-1/1-x is equivalent to? A:none B:1 C:-1 D:x

OpenStudy (anonymous):

\[\left(\begin{matrix}x-1 \\ 1-x\end{matrix}\right)\]

OpenStudy (anonymous):

ok take -1 common from denominator then see what happens !!

OpenStudy (anonymous):

x is left

OpenStudy (anonymous):

let me give you an example if we have \[\Large \frac{x-5}{5-x}\] taking -1 common from denominator \[\Large \frac{(x-5)}{-1(x-5)}\] now since x-5 is in both numerator and denominator they will gets cancelled so. \[\Large \implies \frac{1}{-1}=-1\] do the same for your question !!!!!!

OpenStudy (anonymous):

so it's one

OpenStudy (anonymous):

do it again!!! look at my response again. you are taking -1 common how it can be 1 !!!!

OpenStudy (anonymous):

ok is it -1

OpenStudy (anonymous):

well done !!!!! it is -1 you got it.